My P2 team of Brock, Jesse, Mackenzie, and Tyler challenged my P4 team of Hannah, Kendall, Avery, and Renee to a tug-o-war.
Here is the list of forces during the tug-o-war
P2 team: Brock 450N, Jesse 225N, Mackenzie 200N, Tyler 340N P4 team: Hannah 138N, Kendall 215N (she only used one hand) Avery 274, Renee 138N.
What was the net force? Who won?
1
Expert's answer
2020-04-20T10:12:22-0400
"F_{P2}=450+225+200+340=1215\ N"

"F_{P4}=138+215+274+138=765\ N"

"F_{net}=1215-765=450\ N"



P2 team of Brock, Jesse, Mackenzie, and Tyler wins.